# FFAA: Multimodal Large Language Model based Explainable Open-World Face Forgery Analysis Assistant

### v1

@@ Introduction
Table 3: MIDS deisgn choice ablations (%). We compare the overall generalization performance across the entire OW-FFA-Bench and the model’s robustness across all test sets.
Ablation Study on MIDS.
@@
We first explore the effectiveness of MIDS. Fig. 6 presents the distribution of ’Easy’ and ’Hard’ images determined by the MLLM in different test sets of OW-FFA-Bench, reflecting the difficulty of each test set to some extent. After incorporating MIDS, the accuracy on ’Hard’ images significantly improves, effectively enhancing the model’s robustness. For ’Easy’ images, accuracy remains high, demonstrating that our method effectively distinguishes between easy and difficult samples without any threshold settings. We then ablate several design choices of MIDS in Tab. 3. (1) Mask ’Analysis result’ of answers. Before inputting the answer $\bm{X}_{a}$ into MIDS, we try not to mask the analysis result, which yields 78.2% ACC and is 8.3% lower than the masked scenario. We hypothesize that MIDS tends to take shortcuts by focusing solely on the final result when it is visible. By masking the final result, the model is compelled to learn the relationship between the image and the analysis process. (2) Unfreeze the last two layers of CLIP-ViT. We further unfreeze the last two layers of visual encoder of MIDS and observe further improvements in generalization. However, without masking, there is no improvement in ACC and the model’s robustness declines, indicating that the CLIP-ViT’s ability to perceive face authenticity effectively improves only with attention to enriched textual features from the forgery analysis process.
Qualitative Study
Comparison with Advanced MLLMs.
Fig. 5 depicts a quality example. We present some responses from advanced MLLMs, including LLaVA-v1.6-34b (Liu et al. 2024a) and GPT-4o, and compared them with ours. These advanced MLLMs exhibit strong image understanding capabilities but often struggle to determine face authenticity, sometimes avoiding definitive conclusions. In contrast, our method not only understands the image but also conducts a reasoned analysis of authenticity from various perspectives.
@@
Figure 7: Attention heatmap visualization. The first row sequentially displays two sets of (image, heatmaps for $\bm{F}_{vl}$ and $\bm{F}_{vg}$), while the second row presents key information from FFAA’s answer.
@@
Figure 8: Easy and hard samples visualization. The first and second rows respectively display the ’Easy’ samples and ’Hard’ samples as determined by FFAA on the OW-FFA-Bench.
Heatmap Visualization.
@@
To verify whether MIDS can learn more face authenticity representation features guided by enriched textual features, we utilize Grad-CAM (Selvaraju et al. 2020) to visualize the attention heatmap of $\bm{F}_{vl}$ and $\bm{F}_{vg}$. As illustrated in Fig. 7, guided by enriched textual features, MIDS considers multiple aspects of the image and focuses on key features related to face authenticity.
Visualization of Easy and Hard Samples
@@
As illustrated in Fig. 8, the fake faces in the ’Easy’ samples usually have obvious signs of forgery while real faces have higher image quality. Conversely, ’Hard’ samples often exhibit varying image quality and more convincing forgeries, posing significant challenges in determining face authenticity.
Conclusion
@@
In this paper, we introduce a novel OW-FFA-VQA task aimed at understanding the decision-making process of face forgery analysis models and establish the corresponding benchmark and dataset. To tackle this task, we introduce FFAA: Face Forgery Analysis Assistant, which consists of a fine-tuned MLLM and MIDS. By integrating hypothetical prompts with MIDS, the impact of fuzzy classification boundaries between real and forged faces is effectively mitigated, enhancing the model’s robustness. Our experiments demonstrate that FFAA not only yields user-friendly and explainable results but also significantly boosts generalization, accuracy and robustness compared to previous methods.
Appendix A Appendix
A. Data Sources
@@
FF++ (Original)
Celeb-DF-v2 (Celeb-real, Youtube-real)
DFFD (FFHQ)
@@
FF++ (Deepfakes, FaceSwap)
Celeb-DF-v2 (Celeb-synthesis)
@@
FF++ (Face2Face, NeuralTexture)
DFFD (Faceapp)
@@
DFFD (StyleGAN-FFHQ)
GanDiffFace
@@
Data Source Composition for FFA-VQA.
We collect data from the previously mentioned public datasets. We employ the dlib tool to crop faces from video frames and images, then uniformly resized them to 224x224. Fig. 9a shows some of these face images. The images are categorized by forgery type based on their sources, as shown in Tab. 4. From the dataset, we randomly select 2,000 images, with half designated as the validation set and the other half as the test set. Following the process outlined in Fig. 2b and using the training set of Multi-attack as the image data source, we create face forgery analysis data. This resulted in 20,142 pairs of VQA data, distributed as follows: ’Real’ (8,073), ’Facial attribute manipulation’ (4,014), ’Entire face synthesis’ (4,025), and ’Identity exchange’ (4,030). The average number of tokens in the answers is 203.
Data Source Composition for OW-FFA-Bench.
The data sources for OW-FFA-Bench are shown in Tab. 1. Similarly, we employ the dlib tool to crop faces from video frames or images and uniformly resize them to 224x224. Finally, to balance testing efficiency and effectiveness, we randomly sample 1,000 images from each dataset according to its data distribution. Fig. 9b shows some of the face images from the test sets.
@@
B. Training Details
Fine-tuning Details of the MLLM.
We employ LLaVA-v1.6-mistral-7B (Liu et al. 2024a) as our backbone and fine-tune it with LoRA (rank=32, alpha=48) on the FFA-VQA dataset containing hypothetial prompts using 2 RTX 3090 GPUs, with a learning rate of 1e-4 and a batch size of 16. All other settings remain the same as in LLaVA.
Training Details of MIDS.
We first employ the fine-tuned MLLM to mine historical answer data from the remaining available images in the Multi-attack training set, excluding the ’Easy’ images. Ultimately, we collect 30K $(\bm{X}_{v},\bm{X}_{a}^{a},\bm{X}_{a}^{p},\bm{X}_{a}^{n})$ data pairs. We then train MIDS on this dataset for 2 epochs using 2 RTX 3090 GPUs, with the AdamW optimizer, a learning rate of 1e-4, a batch size of 48, and a weight decay of 1e-5. The number of self-attention iterations $M=3$. Data augmentation techniques include ImageCompression, GaussianNoise, MotionBlur, GaussianBlur, FancyPCA and HueSaturationValue.
Training Details of Baseline.
The training details for the methods we trained on the Multi-attack training set are as follows, with all unmentioned details consistent with the code provided by the original authors:
@@ B. Training Details
C. Answer Analysis
@@
D. Prompts
The system prompt, data source specific prompt, question prompt and reference prompt are shown in Fig. 12.
E. Examples
@@
Figure 12: From top to bottom: system prompt, data source specific prompt, question prompt, reference prompt.
Figure 13: Examples of dialog of FFAA under different lighting conditions. The first image, a forged face with the forgery type of identity exchange, is sourced from DeepFakeDetection (Google 2019), while the second image, showing a real face, is sourced from Deeperforensics (Jiang et al. 2020).
Figure 14: Examples of dialog of FFAA in scenarios involving malicious photo editing and synthetic faces. The first image, a forged face with the forgery type of facial attribute manipulation generated using FaceApp (Limited 2017), while the second image, showing a synthetic face generated using StyleGAN (Karras, Laine, and Aila 2019), both are sourced from DFFD (Dang et al. 2020).
Figure 15: Examples of dialog of FFAA in low-quality image scenarios. The first image, a forged face with the forgery type of identity exchange, is sourced from DFDC (Dolhansky et al. 2020), while the second image, showing a real face, is sourced from Deeperforensics (Jiang et al. 2020).
Figure 16: Examples of dialog of FFAA in scenarios involving obvious forgeries. The first image, a synthetic face generated using StyleGAN (Karras, Laine, and Aila 2019), is sourced from DFFD (Dang et al. 2020), while the second image, a forged face with the forgery type of identity exchange, is sourced from DFDC (Dolhansky et al. 2020).
Figure 17: Comparison of ROC curves for generalization performance tests on the OW-FFA-Bench (single test).
Figure 18: More heatmaps visualization. We present eight face images from different sources, along with the heatmap visualizations of MIDS’s $\bm{F}_{vl}$ and $\bm{F}_{vg}$, and several key information from FFAA’s answers. It can be observed that $\bm{F}_{vl}$ tends to focus on specific facial features, the surrounding environment, and more apparent local forgery marks or authenticity clues, whereas $\bm{F}_{vg}$ emphasizes more abstract features (e.g., ’smoothness’, ’lighting’, ’shadows’, ’integration’). In some cases, such as examples (a), (b) and (c), $\bm{F}_{vl}$ and $\bm{F}_{vg}$ may rely on the same features as clues for assessing the authenticity of the face.

### v2

@@ 5.2 Comparison with Competitive Methods
Table 3: MIDS design choice ablations (%). We compare the overall generalization performance across the entire OW-FFA-Bench and model robustness across all test sets.
5.3 Ablation Study
Figure 6: The distribution of ’Easy’ and ’Hard’ images as determined by the fine-tuned MLLM and the accuracy improvements on ’Hard’ images achieved by MIDS. On ’Easy’ images, performance is consistent both with and without MIDS.
Ablation Study on FFA-VQA: We study the effectiveness of the FFA-CoT and the impact of high-quality image descriptions and forgery reasoning textual data on various MLLMs. Specifically, we utilize advanced and open-source MLLMs, including LLaVA-phi-3-mini [1] and LLaVA-mistral-7B [26]. We fine-tune these MLLMs on FFA-VQA containing only ’Analysis results’ (R), ’Image description + Analysis results’ (DR), ’Forgery reasoning + Analysis results’ (FR) and the complete FFA-CoT (All) respectively. We evaluate the fine-tuned MLLMs on the in-domain test set (MA) and the cross-domain test set (OW-FFA-Bench). Fig. 3 presents the detection accuracy (%) of these MLLMs across various test sets. Relying solely on ’Analysis results’ leads to overfitting and limited generalization. With the inclusion of image descriptions and forgery reasoning, the task evolves into a VQA task that necessitates deeper analysis and reasoning, enhancing generalization on cross-domain test sets. In addition, employing the complete FFA-CoT further improves generalization.
@@ 5.3 Ablation Study
Ablation Study on MIDS: We first explore the effectiveness of MIDS. Fig. 6 presents the distribution of ’Easy’ and ’Hard’ images determined by the MLLM in different test sets of OW-FFA-Bench, reflecting the difficulty of each test set to some extent. After incorporating MIDS, the accuracy on ’Hard’ images significantly improves, effectively enhancing model robustness. For ’Easy’ images, accuracy remains high, demonstrating that our method effectively distinguishes between easy and difficult samples without any threshold settings. We then ablate several design choices of MIDS in Tab. 3. (1) Mask ’Analysis result’ of answers. Before inputting the answer $\mathbf{X}_{a}$ into MIDS, we experiment with leaving the ’Analysis Result’ unmasked, resulting in 78.2% ACC, which is 8.3% lower than the masked scenario. We hypothesize that MIDS may be inclined to take shortcuts by focusing on the visible final result. Masking the final result compels the model to infer the relationship between the image and the analysis process. (2) Unfreeze the last two layers of CLIP-ViT. We unfreeze the last two layers of visual encoder of MIDS and observe further improvements in generalization. However, without masking, there is no improvement in ACC and model robustness declines, indicating that the CLIP-ViT’s ability to perceive face authenticity effectively improves only with attention to enriched textual features from the forgery analysis process.
5.4 Qualitative Study
Comparison with Advanced MLLMs in Real-World Scenarios: Fig. 5 illustrates an application scenario in public information security. We compare the responses of FFAA with other advanced MLLMs, including LLaVA-Llama-8B [27] and GPT-4o [32]. LLaVA demonstrates discrimination capabilities but lacks substantial evidential support. GPT-4o exhibits comprehensive knowledge in face forgery analysis, employing multi-perspective assessments but often avoids definitive judgments, likely due to the inherent ambiguity of forgery analysis and safety policies of GPT. FFAA performs a multi-faceted analysis of facial features, extracting detailed forgery clues, assessing authenticity, and providing a match score and difficulty type as indicators of response reliability.
@@ 5.4 Qualitative Study
Figure 7: Attention heatmap visualization. The first row displays two sets of (image, heatmaps for $\mathbf{F}_{vl}$ and $\mathbf{F}_{vg}$), while the second row presents some key information from FFAA’s answer.
@@
Figure 8: Easy and hard samples visualization. The first and second rows respectively display the ’Easy’ samples and ’Hard’ samples as determined by FFAA on OW-FFA-Bench.
@@
Heatmap Visualization: To verify whether MIDS can learn more face authenticity representation features guided by enriched textual features, we utilize Grad-CAM [37] to visualize the attention heatmap of $\mathbf{F}_{vl}$ and $\mathbf{F}_{vg}$. As illustrated in Fig. 7, guided by enriched textual features, MIDS considers multiple aspects of the image and focuses on key features related to face authenticity.
@@
Visualization of Easy and Hard Samples: As illustrated in Fig. 8, the fake faces in the ’Easy’ samples usually have obvious signs of forgery while real faces have higher image quality. Conversely, ’Hard’ samples often exhibit varying image quality and more convincing forgeries, posing significant challenges in determining face authenticity.
6 Conclusion
@@
In this paper, we introduce a novel OW-FFA-VQA task aimed at understanding the decision-making process of face forgery analysis models and establish a corresponding benchmark. We first leverage GPT4-assisted analysis generation to create the FFA-VQA dataset, encompassing diverse real and forged face images with essential descriptions and forgery reasoning. Based on this dataset, we introduce FFAA: Face Forgery Analysis Assistant, which consists of a fine-tuned MLLM and MIDS. Our experiments demonstrate that FFAA not only provides user-friendly and explainable results but also significantly enhances generalization, accuracy and robustness compared to previous methods.
Limitations: Our method prioritizes detailed analysis and model decision transparency, leading to longer inference times and limited suitability for large-scale forgery screening. Additionally, it is restricted to face images, whereas forgery can involve multiple modalities.
Future Works: We will work on reducing inference time while improving accuracy and explore explainable face forgery analysis methods with multi-modal inputs using Multimodal Large Language Models.
Supplementary Material
7 Dataset Construction
Algorithm 1 FFA-VQA Dataset Construction
1:source dataset $\mathcal{M}$, question prompt set $\mathcal{O}$, queried images set $\mathcal{Q}$, system prompt $S$, threshold $\delta$, budget $N$
2:$(\mathcal{I},\mathcal{R})\leftarrow\text{Decompose}(\mathcal{M})$
3:Remaining images set $\mathcal{T}\leftarrow\mathcal{I}\setminus\mathcal{Q}$
4:$k\leftarrow\min(N,|\mathcal{T}|)$
5:Candidate set $\mathcal{C}\leftarrow\emptyset$
6:repeat
7:    Randomly select a question prompt $x_{q}\in\mathcal{O}$
8:    Sample an image $I$ from dataset $\mathcal{T}$
9:    $(L,F)\leftarrow\text{ParseImage}(I)$triangleright $L$: label, $F$: forgery type
10:    $x_{f}\leftarrow\text{GetForgerySpecificPrompt}(L,F)$
11:    if $F$ is ’IE’ or ’FAM’ then
12:         $(R,x_{r})\leftarrow\text{GetReferenceAndPrompt}(I,\mathcal{R})$
13:         $x\leftarrow(I,x_{f},R,x_{r},x_{q})$
14:    else
15:         $R\leftarrow\text{None}$
16:         $x\leftarrow(I,x_{f},x_{q})$
17:    end if
18:    Get answer $A\leftarrow\text{GPT4o}(S,x)$
19:    $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{I\},\;\mathcal{T}\leftarrow\mathcal{T}\setminus\{I\},\;k\leftarrow k-1$
20:    if $A$ does not follow FFA-CoT format then
21:         continue
22:    end if
23:    $(L_{a},F_{a},P_{a})\leftarrow\text{ParseAnswer}(A)$
24:    if $(L_{a}=L)$ and $(F_{a}=F)$ and $(P_{a}\geq\delta)$ then
25:         $\mathcal{C}\leftarrow\mathcal{C}\cup\{(I,R,L,F,x_{q},A)\}$
26:    else
27:         continue
28:    end if
29:until $k\leq 0$
30:$\mathcal{C}^{\prime}\leftarrow\emptyset$
31:for each $(I,R,L,F,x_{q},A)\in\mathcal{C}$ do
32:    if Expert approves $A$ based on $(I,R,L,F)$ then
33:         $\mathcal{C}^{\prime}\leftarrow\mathcal{C}^{\prime}\cup\{(I,x_{q},A)\}$
34:    end if
35:end for
36:Return Qualified VQA dataset $\mathcal{C}^{\prime}$
@@ 7 Dataset Construction
FF++[36] (Original)
Celeb-DF-v2[22] (Celeb-real, Youtube-real)
DFFD[7] (FFHQ)
@@
FF++[36] (Deepfakes, FaceSwap)
Celeb-DF-v2[22] (Celeb-synthesis)
@@
FF++[36] (Face2Face, NeuralTexture)
DFFD[7] (Faceapp)
@@
DFFD[7] (StyleGAN-FFHQ)
GanDiffFace[30]
@@
Data Source Composition for FFA-VQA: We leverage data gathered from the public datasets illustrated in Fig. 2a. Using the dlib library, we extract facial regions from video frames and images, subsequently resizing them uniformly to 224×224. Fig. 9a presents examples of these processed face images. The images are classified by forgery type based on their respective sources, as detailed in Tab. 4. From this dataset, we randomly select 2,000 images, dividing them equally into a validation set and a test set. Following the process outlined in Algorithm 1 and Fig. 2b, and using the training set of Multi-attack as the image data source, we construct the FFA-VQA dataset, comprising 20,142 pairs of VQA data distributed as follows: ’Real’ (8,073), ’Facial attribute manipulation’ (4,014), ’Entire face synthesis’ (4,025), and ’Identity exchange’ (4,030). To balance the difficulty and reliability of face forgery analysis, we set the threshold $\delta$ to 0.6. The average number of tokens in the answers is 203.
Data Source Composition for OW-FFA-Bench: The data sources for OW-FFA-Bench are shown in Tab. 2. Similarly, we employ the dlib tool to crop faces from video frames or images and uniformly resize them to 224x224. Finally, to balance testing efficiency and effectiveness, we randomly sample 1,000 images from each dataset according to its data distribution. Fig. 9b shows some of the face images from the test sets.
8 Implementation Details
Fine-tuning Details of the MLLM: We employ LLaVA-v1.6-mistral-7B [26] as our backbone and fine-tune it with LoRA (rank=32, alpha=48) on the FFA-VQA dataset containing hypothetial prompts using 2 RTX 3090 GPUs, with a learning rate of 1e-4 and a batch size of 16. All other settings remain the same as in LLaVA.
Training Details of MIDS: We first employ the fine-tuned MLLM to mine historical answer data from the remaining available images in the Multi-attack training set, excluding the ’Easy’ images. Ultimately, we collect 30K $(\mathbf{X}_{v},\mathbf{X}_{a}^{a},\mathbf{X}_{a}^{p},\mathbf{X}_{a}^{n})$ data pairs. We then train MIDS on this dataset for 2 epochs using 2 RTX 3090 GPUs, with the AdamW optimizer, a learning rate of 1e-4, a batch size of 48, and a weight decay of 1e-5. The number of self-attention iterations $M=3$. Data augmentation techniques include ImageCompression, GaussianNoise, MotionBlur, GaussianBlur, FancyPCA and HueSaturationValue.
Training Details of Baseline: The training details for the methods we trained on the Multi-attack training set are as follows, with all unmentioned details consistent with the code provided by the original authors:
  • ResNet-34: We trained for 20 epochs on 2 RTX 3090 GPUs using the Adam optimizer, with a learning rate of 1e-3, a batch size of 100, and a weight decay of 1e-6. Data augmentation included HorizontalFlip.
  • EfficientNet-B4: We trained for 20 epochs on 2 RTX 3090 GPUs using the Adam optimizer, with a learning rate of 1e-3, a batch size of 128, and a weight decay of 1e-6. Data augmentation included HorizontalFlip.
@@ 8 Implementation Details
  • DCL: We trained for 30 epochs on 2 RTX 3090 GPUs using the Adam optimizer, with a learning rate of 2e-4, a batch size of 32, and a weight decay of 1e-5. Additionally, the parameter has_mask was set to false.
@@
  • UIA-ViT: We trained for 8 epochs on 1 RTX 3090 GPU using the Adam optimizer, with a learning rate of 3e-5, a batch size of 96.
@@
  • F3Net: We trained for 6 epochs on 1 RTX 3090 GPU using the Adam optimizer, with a learning rate of 2e-4, a batch size of 128.
  • Two-Stream: We trained for 5 epochs on 2 RTX 3090 GPUs using the Adam optimizer, with a learning rate of 2e-4, a batch size of 64, and a weight decay of 1e-6. Data augmentation included HorizontalFlip.
  • FADD: We trained for 40 epochs on 1 RTX 3090 GPUs using the Adam optimizer, with a learning rate of 1e-3, a batch size of 64.
9 Answer Analysis
@@
10 Prompts
The system prompt, forgery-specific prompt, question prompt and reference prompt are shown in Fig. 12.
11 Examples
Examples of FFA-VQA dataset: Fig. 11 illustrates examples of VQA pairs for four categories of faces in the FFA-VQA dataset. It can be observed that the image descriptions comprehensively detail facial features, image quality, and surrounding environments. Additionally, the forgery reasoning section provides robust evidence for analysis.
Dialogue Examples of FFAA: Figures 13 to 16 present FFAA dialogue examples from the OW-FFA-Bench and Multi-attack test set, covering scenarios such as varied lighting, image editing, generated images, low resolution, and evident forgery traces. Figures 17 to 20 highlight FFAA’s performance against real-world unknown forgery techniques, including celebrity face swaps and high-fidelity AIGC-generated faces. These examples demonstrate FFAA’s ability to provide strong evidence and accurately identify forgery types, even when faced with unknown deepfake technologies.
More Heatmaps Visualization: Fig. 21 presents additional heatmap visualization results, further illustrating how MIDS, guided by rich semantic analysis, effectively captures detailed image features and diverse, deeper forgery characteristics.
@@ 11 Examples
Figure 12: From top to bottom: system prompt, forgery-specific prompt, question prompt, reference prompt.
Figure 13: Dialogue examples of FFAA under different lighting conditions. The first image, a forged face with the forgery type of identity exchange, is sourced from DeepFakeDetection [11], while the second image, showing a real face, is sourced from Deeperforensics [15].
Figure 14: Dialogue examples of FFAA in scenarios involving malicious photo editing and synthetic faces. The first image, a forged face with the forgery type of facial attribute manipulation generated using FaceApp [23], while the second image, showing a synthetic face generated using StyleGAN [17], both are sourced from DFFD [7].
Figure 15: Dialogue examples of FFAA in low-quality image scenarios. The first image, a forged face with the forgery type of identity exchange, is sourced from DFDC [8], while the second image, showing a real face, is sourced from Deeperforensics [15].
Figure 16: Dialogue examples of FFAA in scenarios involving obvious forgeries. The first image, a synthetic face generated using StyleGAN [17], is sourced from DFFD [7], while the second image, a forged face with the forgery type of identity exchange, is sourced from DFDC [8].
Figure 17: Dialogue examples of FFAA in real-world scenarios involving celebrity face-swapping. The first image features Elon Musk’s face sourced from the internet, while the second image is a forged face replaced using an unknown forgery technique by another person.
Figure 18: Dialogue examples of FFAA in real-world scenarios involving celebrity face-swapping. The first image features a real face sourced from the internet, while the second image is a forged face replaced using an unknown forgery technique by another person.
Figure 19: Dialogue examples of FFAA in real-world scenarios involving celebrity face-swapping. The first image is a facial screenshot of Leonardo DiCaprio from a movie, sourced from the internet, while the second image is a forged face replaced using an unknown forgery technique by another person.
Figure 20: Dialogue examples of FFAA in real-world scenarios involving AIGC-generated faces. Both images are high-fidelity depictions of beauty, sourced from the internet and created using unknown generation techniques.
Figure 21: More heatmaps visualization. We present sixteen face images from different sources, along with the heatmap visualizations of MIDS’s $\mathbf{F}_{vl}$ and $\mathbf{F}_{vg}$, and several key information from FFAA’s answers. It can be observed that $\mathbf{F}_{vl}$ tends to focus on specific facial features, the surrounding environment, and more apparent local forgery marks or authenticity clues, whereas $\mathbf{F}_{vg}$ emphasizes more abstract features (e.g., ’smoothness’, ’lighting’, ’shadows’, ’integration’). In some cases, such as examples (a), (b), (c), (l), (m), (o) and (p), $\mathbf{F}_{vl}$ and $\mathbf{F}_{vg}$ may rely on the same features as clues for assessing the authenticity of the face.